Test of DB#F4 on MIR 556 with 10 slews. All ok.

Work starts with substitution of DB#F6 with DB#F5 removed form the system and giving issues on MIR#556 in order to fix the problems related to DB#F5 (managing MIR450)

- Reading OK in MIR528 but 180u, instead of 160um read and drifting up. (Has been substituted with spare in Arcetri, so probably cannot be used in closed loop.).

- Removed flat cable on MIR611 ( reading now on MIR528 from 180 drift up, as before---> some other problem....)

- Test on MIR450 (was jumping). The signal is not visible (is 2m). Releasing the DB and pushing it towards the center of the unit, the signal appears.

Found one siggen with HALF signal (giving BAD readings(double) on capses) color YELLOW -> crate 1.

Even testing actuators on other crate on that siggen the signal is still HALF.

- repeated test on act MIR528 with a good siggen: not drifting

- repeated test on act MIR450 with a good siggen: same as before, visible only moving the DB.

The SGN serial numbers are not coherent with MG report: read SN (11, 5, 12, 6, 13, 7) and MG reported substitution 5 with 11 (the 5 is present but on crate 1)

The MIR450 was substituted with a spare actuator (SN727 with SN706)

- test on MIR450: ALL OK after 20 slews---> it means that probably the problems on MIR556 was not related to DB but to the coupling between actuator and DB connector (maybe due to mechanics?)

- reinserted flat cable on MIR611 in order to verify that actuator with a good siggen board: OPEN , still not working

The MIR611 has coil lines in contact to cap sensor armature. Substituted with spare one (SN1671 with SN1680 (only one found good in coil lines))

All others external spares have same problem on coil lines (see picture attached.)

The one substituted was previously replaced but now seems to go well... (10 slew tests)--> was written MIR651.

Test on 379MIR 5 slew test: no problems.

Test slew on MIR317: no problems with slew but identifies with telemetry as critical (vulcano effect). Substituted (SN1069 with SN707)

Substituted actuators MIR471, MIR472, MIR474 (SN980 with SN708, SN1020 with SN709, SN1005 with SN710)

Changed accelerometer crate 1 with spare (SN5 with SN8). Tested with ac couple of actuators (crate 0 and crate 1) and all is ok.

MIR424 went 2m, MIR390 half signal duriing tests. Subtituted: (390:SN1207 with SN711) and (424:SN1245 it was decided to remove the capsens for the difficult due to the insertion, leaving only the cold finger. The insertion of cold finger remained difficult so the clamp is the reason. We substituted at the end SN1245 with SN713).

MIR589(DSP31) external actuator removed for inspection for the shortcut of coils wires inside RB hole.It was touching BUT not causing problems thanks to a large glue spot but it is touching. So we change it with a spare (SN1650 with 1681)

Quick test with 3< slews each for verifying no damage: MIR471, MIR472, MIR474, MIR317, MIR390, MIR589, MIR424 (all ok)

Shell mounted. No contamination found.

Actuators DSP281/MIR320 started with 2m signal. Never been noticed before. Removed from good actuators.

Shell set for membrane adjustment. Found 4 external actuators on DB#F6 not working properly (DSP340-341-342-343). Large drift detected.

Was demonstrated that MIR556 problem was not related to DB (see log above). Therefore we proceed with back substitution DB#F6 with DB#F5 (remove spare DB#F2 ----> FOUND FAULTY) *

Found a log mismatch. the full series of changes of DB is:

original F6 on MIR556, F5 on MIR450

F2 replaced F6, and F5 stay there.

F2 replaced F6 and F6 replaced F5.

F5 replaced F6.

At this moment we recovered actuators: MIR450, MIR556, MIR611, MIR379, MIR471, MIR472, MIR474, MIR317, MIR390, MIR424, MIR589


The actuators MIR319/DSP280 has a small drift. Removed from ICC.

Found little vulcano effect: Wed Aug 24 22:47:09 GMT 2011 (telemetry) but does not seem relevant (around 320/329MIR??? TBC) -> 3e-3 forces with only FF gain and ICC

10 minutes slews departing: Wed Aug 24 22:51:37 GMT 2011

confirmed 568,338, 315, 327

Shell RIP Wed Aug 24 22:57:03 GMT 2011 ---> probably forgot ELEVATION setting

Started again: Wed Aug 24 23:00:55 GMT 2011 -> Wed Aug 24 23:11:00 GMT 2011

Slew stopping every 10 deg: Wed Aug 24 23:11:46 GMT 2011 -> Wed Aug 24 23:16:28 GMT 2011

Analyzed FF residuals: TN 20110824_231832 (zonal)

Decide to remove actuator 280.

-- MarcoXompero - 24 Aug 2011
  • drift.png:
    drift.png

  • orecchia.png:
    orecchia.png

  • residue.png:
    residue.png
Topic attachments
I Attachment Action Size Date Who Comment
drift.pngpng drift.png manage 8 K 25 Aug 2011 - 02:58 MarcoXompero  
orecchia.pngpng orecchia.png manage 54 K 25 Aug 2011 - 02:58 MarcoXompero  
residue.pngpng residue.png manage 23 K 25 Aug 2011 - 05:53 MarcoXompero  
Topic revision: r5 - 25 Aug 2011, MarcoXompero
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback